# DisTaC: Conditioning Task Vectors via Distillation for Robust Model Merging

### v1

@@ Abstract
Model merging has emerged as an efficient and flexible paradigm for multi-task learning, with numerous methods being proposed in recent years. However, these state-of-the-art techniques are typically evaluated on benchmark suites that are highly favorable to model merging, and their robustness in more realistic settings remains largely unexplored. In this work, we first investigate the vulnerabilities of model-merging methods and pinpoint the source-model characteristics that critically underlie them. Specifically, we identify two factors that are particularly harmful to the merging process: (1) disparities in task vector norms, and (2) the low confidence of the source models. To address this issue, we propose DisTaC (Distillation for Task vector Conditioning), a novel method that pre-conditions these problematic task vectors before the merge. DisTaC leverages knowledge distillation to adjust a task vector’s norm and increase source-model confidence while preserving its essential task-specific knowledge. Our extensive experiments demonstrate that by pre-conditioning task vectors with DisTaC, state-of-the-art merging techniques can successfully integrate models exhibiting the harmful traits—where they would otherwise fail—achieving significant performance gains. The source code is available at https://github.com/katoro8989/DisTaC
@@ 1 Introduction
The recent wave of open-sourcing both large pretrained models (Devlin et al., 2019; Rombach et al., 2022; Achiam et al., 2023; Grattafiori et al., 2024) and their fine-tuned downstream variants (Wolf et al., 2019; Taori et al., 2023) has put an unprecedented variety of neural networks within easy reach of anyone. This democratization has, in turn, accelerated research on model merging (Wortsman et al., 2022b; a; Ilharco et al., 2023; Yadav et al., 2023; Akiba et al., 2025)—techniques that create new, customized models by integrating existing fine-tuned models without the need for additional large-scale training. In particular, a flurry of recent methods aims to build multi-task models by merging networks that have been fine-tuned independently for each task, rather than retraining a single shared model from scratch (Ilharco et al., 2023; Yadav et al., 2023; Ortiz-Jimenez et al., 2023; Wang et al., 2024; Yoshida et al., 2025; Gargiulo et al., 2025). Many of these techniques require only minimal extra training—or none at all. Compared with conventional multi-task learning (MTL), they offer two key advantages: (i) they eliminate the need to aggregate all task-specific labeled data in one location, sidestepping data-access constraints, and (ii) they make it easy to add or edit the model’s skill on a particular task after deployment (Yang et al., 2024a).
@@
On established benchmarks, these approaches have shown promising gains, in some cases approaching the performance of traditional MTL (Gargiulo et al., 2025). Yet those benchmarks are built under conditions that are highly idealized for model merging; how robust current merging methods remain in more practical, pessimistic settings is still largely unknown. Bridging this gap is a prerequisite for real-world application.
@@
To that end, we first pinpointed where generic multi-task model merging pipelines break down. Our analysis reveals two especially harmful factors: (1) differences in task vector norms and (2) low prediction confidence of source models. Figure 1(a) illustrates the vulnerability of recent merging methods to these factors using CLIP (Radford et al., 2021) with a ViT-B-32 (Dosovitskiy et al., 2021) backbone on the eight vision tasks defined in Section 5.1: blue bars show the effect of training models with different learning rates, thereby altering task vector norms (see Figure 1(b)), while yellow bars show the effect of label smoothing (Müller et al., 2019) (LS), which reduces model confidence (see Figure 1(c)). In the plot, the horizontal axis lists the merging methods, and the vertical axis reports the average normalized accuracy (Norm. ACC) across the eight tasks, defined as the post-merge accuracy relative to the pre-merge accuracy obtained by individual models for each task. In both cases, every method’s performance degrades substantially compared to the standard baseline, represented by the gray bars (a uniform learning rate of $10^{-5}$ with hard labels), with a maximum 24% drop in Norm. ACC.
@@
These failure modes often arise in real-world deployments. For instance, differences in task vector norms can stem from varied learning rates, fine-tuning steps, or weight decay used during the individual fine-tuning of each task (Devlin et al., 2019; Wightman et al., 2021). Low confidence often results from techniques such as LS, Mixup (Zhang et al., 2017), and focal loss (Lin et al., 2017). We therefore contend that models should be pre-conditioned before merging to remove their latent harmfulness. To this end, we propose DisTaCDistillation for Task-vector Conditioning—a lightweight knowledge distillation (KD) procedure that tackles both issues using only unlabeled data: To correct task vector–norm disparities, DisTaC first rescales each vector to a chosen target norm and then restores any performance lost through this scaling by distilling knowledge from the original model. To address low source-model confidence, it trains the student with a higher temperature than the teacher ($T_{\text{stu}}>T_{\text{tcr}}$), so the student ultimately produces lower-entropy—that is, more confident—predictions. Algorithm 1 combines these two conditioning steps, allowing them to be carried out in a single pass. Because DisTaC leverages the already-trained task vectors as the initialization for KD and relies solely on unlabeled data, it incurs minimal computational overhead and imposes only modest practical requirements, yet markedly improves the robustness of existing model merging techniques in challenging scenarios.
Empirically, on eight vision tasks with ViT-B-32/L-14 backbones, DisTaC increased post-merge accuracy by up to 20.8 percentage points and restored the best-performing TSVM merge’s normalized accuracy from 68% to 92% under low-confidence conditions, thereby matching the conventional “ideal” benchmark performance (i.e., merging high-confidence models with uniform task vector norms)—all with minimal computational cost.
Our contributions are as follows:
@@
  • From our analysis, we present two practical guidelines for practical model merging: (i) when the task vector norms differ, it is better to shrink the larger vector rather than stretch the smaller one (Section 6.1); and (ii) when the source models have low confidence, it is more effective to make them overconfident before merging, and then apply a calibration method to the merged model (Section 6.2).
@@ 2 Preliminaries
For any parameter vector $\bm{\theta}\in\mathbb{R}^{d}$, let $f(\cdot\,;\bm{\theta}):\mathcal{X}\to\mathbb{R}^{C}$ denote the neural network that maps an input $x\!\in\!\mathcal{X}$ to a $C$-dimensional logit vector, where $C$ is the number of classes. We target a multi-task scenario comprising $T$ supervised tasks. Let $\bm{\theta}_{\text{pre}}\in\mathbb{R}^{d}$ be the parameters of an open-source pretrained backbone. For each task $t\in\{1,\dots,T\}$, we obtain a model that has already been fine-tuned on the corresponding labelled dataset $\mathcal{D}_{t}=\bigl{\{}(x_{i}^{(t)},y_{i}^{(t)})\bigr{\}}_{i=1}^{n_{t}},$ yielding task-specific weights $\bm{\theta}_{t}\in\mathbb{R}^{d}$.
@@ 2.1 Model Merging for Multi-Task Learning
Recent model merging techniques operate on the task vectors (Ilharco et al., 2023) $\bm{\tau}_{t}\;:=\;\bm{\theta}_{t}-\bm{\theta}_{\text{pre}}$ and obtain a single multi–task model by linearly combining them:
$\bm{\theta}_{\text{mtl}}\;=\;\bm{\theta}_{\text{pre}}\;+\;\sum_{t=1}^{T}P_{t}\,\bm{\tau}_{t},$ (1)
where each $P_{t}\in\mathbb{R}^{d\times d}$ is a method-specific matrix that mitigates inter-task interference.
In the following, we explain the $P_{t}$ used in each merging method.
Uniform averaging: $P_{t}=\tfrac{1}{T}\,I_{d}$.
Task arithmetic (Ilharco et al., 2023): $P_{t}=\lambda_{t}\,I_{d},\quad\text{where}\ \lambda_{t}\in\mathbb{R}$.
Ties-Merging (Yadav et al., 2023): $P_{t}=\lambda_{t}\,\bm{m}_{\text{Ties},t}\,I_{d},\quad\text{where}\ \lambda_{t}\in\mathbb{R},\ \bm{m}_{\text{Ties},t}\in\{0,1\}^{d}$. $\bm{m}_{\text{Ties},t}$ is determined by the norm of each weight parameter to mitigate inter-task conflicts.
Consensus Merging (Wang et al., 2024): $P_{t}=\lambda_{t}\,\bm{m}_{\text{Cons},t}\,I_{d},\quad\text{where}\ \lambda_{t}\in\mathbb{R},\ \bm{m}_{\text{Cons},t}\in\{0,1\}^{d}$. The framework is the same as Ties-Merging, but the binary mask $\bm{m}_{\text{Cons},t}$ is determined in the following steps. First, create the TALL mask $\bm{m}_{\text{TALL},t}$, which is a binary mask of weights where each element is set to 1 if the norm of $\bm{\tau}_{t}$ is larger than the weighted distance between $\bm{\tau}_{t}$ and $\sum_{t=1}^{T}\bm{\tau}_{t}$. Then, create $\bm{m}_{\text{Cons},t}$, where each element is set to 1 if the corresponding element of $\bm{m}_{\text{TALL},t}$ is 1 in at least $k$ tasks, reflecting agreement among the source models regarding the importance.
TSVM (Gargiulo et al., 2025) cannot be expressed within the framework of Eq. (1). Instead, it suppresses task interference by whitening the matrices $\mathbf{U}_{t}$ and $\mathbf{V}_{t}$ obtained from the singular value decomposition of the task vectors $\bm{\tau}_{t}=\mathbf{U}_{t}\mathbf{\Sigma}_{t}\mathbf{V}_{t}^{\top}$.
@@ 2.2 Knowledge Distillation
Knowledge distillation (KD) is a model compression and transfer paradigm in which a compact student network is trained to replicate the behavior of a larger, well-performing teacher network (Hinton et al., 2015). By minimizing a joint loss that combines ground-truth supervision with a soft-target signal derived from the teacher’s output distribution, the student acquires the teacher’s dark knowledge—fine-grained inter-class relationships encoded in the soft logits—while retaining a substantially smaller parameter footprint. Formally, for a given input $\bm{x}$, let $\bm{z}_{\text{tcr}}:=f(\bm{x}\,;\bm{\theta}_{\text{tcr}})\in\mathbb{R}^{C}$ and $\bm{z}_{\text{stu}}:=f(\bm{x}\,;\bm{\theta}_{\text{stu}})\in\mathbb{R}^{C}$ be the output logits from the teacher and student models, parameterized by $\bm{\theta}_{\text{tcr}}\in\mathbb{R}^{d}$ and $\bm{\theta}_{\text{stu}}\in\mathbb{R}^{d}$, respectively. The KD objective then augments the conventional cross-entropy loss $\mathcal{L}_{\text{CE}}$ with a softened Kullback-Leibler (KL) divergence term:
$\mathcal{L}_{\text{KD}}=(1-\zeta)\,\mathcal{L}_{\text{CE}}\bigl{(}\bm{z}_{\text{stu}},\,y\bigr{)}\;+\;\zeta\ T_{\text{tcr}}T_{\text{stu}}\,\mathrm{KL}\Bigl{(}\sigma(\bm{z}_{\text{tcr}}/T_{\text{tcr}})\,\bigl{\|}\,\sigma(\bm{z}_{\text{stu}}/T_{\text{stu}})\Bigr{)},$ (2)
@@ 3.1 Task Vector Norm Disparity
To quantify this effect, we fine-tuned CLIP models with Vision Transformers (ViTs) backbones, specifically ViT-B-32, on eight vision tasks as introduced in Section 5.1 with two learning rates, $10^{-5}$ (gray) and $10^{-4}$ (blue), and plotted the resulting task vector norms in Figure 1(b). Across all tasks, we observe a 555-7×7$\times$7 × gap between the two settings. Crucially, the difference is not confined to any particular layer: parameter scales diverge consistently throughout the network, as demonstrated in Section C.1.
[…]
The detrimental effect of norm disparity on model merging can be explained with a straightforward theoretical analysis. Let us consider two task vectors, $\bm{\tau}_{1}$ and $\bm{\tau}_{2}$, derived from different tasks, and assume their norms have a large discrepancy, i.e., $||\bm{\tau}_{2}||\gg||\bm{\tau}_{1}||$. In a simple addition merge, the resulting task vector is $\bm{\tau}_{\text{merge}}=\bm{\tau}_{1}+\bm{\tau}_{2}$. The directional alignment of this merged vector can be analyzed using cosine similarity.
The similarity between $\bm{\tau}_{\text{merge}}$ and the high-norm vector $\bm{\tau}_{2}$ approaches one. By expanding the cosine similarity definition, we have $\cos(\bm{\tau}_{\text{merge}},\bm{\tau}_{2})=(\bm{\tau}_{1}\cdot\bm{\tau}_{2}+||\bm{\tau}_{2}||^{2})/(||\bm{\tau}_{1}+\bm{\tau}_{2}||\cdot||\bm{\tau}_{2}||)$. Since $||\bm{\tau}_{2}||\gg||\bm{\tau}_{1}||$, the numerator is dominated by the $||\bm{\tau}_{2}||^{2}$ term, while in the denominator, $||\bm{\tau}_{1}+\bm{\tau}_{2}||\approx||\bm{\tau}_{2}||$. This leads to the approximation
$\cos(\bm{\tau}_{\text{merge}},\bm{\tau}_{2})\approx||\bm{\tau}_{2}||^{2}/(||\bm{\tau}_{2}||\cdot||\bm{\tau}_{2}||)=1.$
Conversely, its alignment with the low-norm vector $\bm{\tau}_{1}$ becomes negligible. The corresponding similarity is $\cos(\bm{\tau}_{\text{merge}},\bm{\tau}_{1})=(||\bm{\tau}_{1}||^{2}+\bm{\tau}_{1}\cdot\bm{\tau}_{2})/(||\bm{\tau}_{1}+\bm{\tau}_{2}||\cdot||\bm{\tau}_{1}||)$. Assuming the tasks are orthogonal ($\bm{\tau}_{1}\cdot\bm{\tau}_{2}=0$), as observed by Ilharco et al. (2023), and using the same approximation for the denominator, the expression simplifies to
$\cos(\bm{\tau}_{\text{merge}},\bm{\tau}_{1})\approx||\bm{\tau}_{1}||^{2}/(||\bm{\tau}_{2}||\cdot||\bm{\tau}_{1}||)=||\bm{\tau}_{1}||/||\bm{\tau}_{2}||\approx 0.$
This demonstrates that the merged solution almost entirely inherits the directional characteristics of the high-norm task, leading to a catastrophic drop in performance for the task associated with the low-norm vector and thus degrading the overall utility of the merged model.

### v2

# arXiv:2508.01148v2[cs.LG] 27 Feb 2026

@@ Abstract
Model merging has emerged as an efficient and flexible paradigm for multi-task learning, with numerous methods being proposed in recent years. However, these state-of-the-art techniques are typically evaluated on benchmark suites that are highly favorable to model merging, and their robustness in more realistic settings remains largely unexplored. In this work, we first investigate the vulnerabilities of model-merging methods and pinpoint the source-model characteristics that critically underlie them. Specifically, we identify two factors that are particularly harmful to the merging process: (1) disparities in task vector norms, and (2) the low confidence of the source models. To address this issue, we propose DisTaC (Distillation for Task vector Conditioning), a novel method that pre-conditions these problematic task vectors before the merge. DisTaC leverages knowledge distillation to adjust a task vector’s norm and increase source-model confidence while preserving its essential task-specific knowledge. Our extensive experiments demonstrate that by pre-conditioning task vectors with DisTaC, state-of-the-art merging techniques can successfully integrate models that exhibit these harmful traits, where they would otherwise fail, and achieve significant performance gains. The source code is available at https://github.com/katoro8989/DisTaC
@@ 1 Introduction
The recent wave of open-sourcing both large pretrained models (devlin2019bert; rombach2022high; achiam2023gpt; grattafiori2024llama) and their fine-tuned downstream variants (wolf2019huggingface; taori2023alpaca) has put an unprecedented variety of neural networks within easy reach of anyone. This democratization has, in turn, accelerated research on model merging (wortsman2022robust; wortsman2022soups; ilharco2023editingmodelstaskarithmetic; yadav2023tiesmerging; akiba2025evolutionary), techniques that create new, customized models by integrating existing fine-tuned models without the need for additional large-scale training. In particular, a flurry of recent methods aims to build multi-task models by merging networks that have been fine-tuned independently for each task, rather than retraining a single shared model from scratch (ilharco2023editingmodelstaskarithmetic; yadav2023tiesmerging; ortiz-jimenez2023task; wang2024localizing; yoshida2025mastering; gargiulo2025task). Many of these techniques require only minimal extra training or none at all. Compared with conventional multi-task learning (MTL), they offer two key advantages: (i) they eliminate the need to aggregate all task-specific labeled data in one location, sidestepping data-access constraints, and (ii) they make it easy to add or edit the model’s skill on a particular task after deployment (yang2024model).
@@
On established benchmarks, these approaches have shown promising gains, in some cases approaching the performance of traditional MTL (gargiulo2025task). Yet those benchmarks are built under conditions that are highly idealized for model merging; how robust current merging methods remain in more practical, pessimistic settings is still largely unknown. Bridging this gap is a prerequisite for real-world application.
@@
To that end, we first pinpointed where generic multi-task model merging pipelines break down. Our analysis reveals two especially harmful factors: (1) differences in task vector norms and (2) low prediction confidence of source models. Figure 1(a) illustrates the vulnerability of recent merging methods to these factors using CLIP (radford2021learning) with a ViT-B-32 (dosovitskiy2021an) backbone on the eight vision tasks defined in Section 5.1: blue bars show the effect of training models with different learning rates, thereby altering task vector norms (see Figure 1(b)), while yellow bars show the effect of label smoothing (muller2019does) (LS), which reduces model confidence (see Figure 1(c)). In the plot, the horizontal axis lists the merging methods, and the vertical axis reports the average normalized accuracy (Norm. ACC) across the eight tasks, defined as the post-merge accuracy relative to the pre-merge accuracy obtained by individual models for each task. In both cases, every method’s performance degrades substantially compared to the standard baseline, represented by the gray bars (a uniform learning rate of $10^{-5}$ with hard labels), with a maximum 24% drop in Norm. ACC.
@@
These failure modes often arise in real-world deployments. For instance, differences in task vector norms can stem from varied learning rates, fine-tuning steps, or weight decay used during the individual fine-tuning of each task (devlin2019bert; wightman2021resnet). Low confidence often results from techniques such as LS, Mixup (zhang2017mixup), and focal loss (lin2017focal). We therefore contend that models should be pre-conditioned before merging to remove their latent harmfulness. To this end, we propose Distillation for Task-vector Conditioning (DisTaC) a lightweight knowledge distillation (KD) procedure that tackles both issues using only unlabeled data: To correct task vector–norm disparities, DisTaC first rescales each vector to a chosen target norm and then restores any performance lost through this scaling by distilling knowledge from the original model. To address low source-model confidence, it trains the student with a higher temperature than the teacher ($T_{\text{stu}}>T_{\text{tcr}}$), so the student ultimately produces lower-entropy predictions, that is, predictions that are more confident.
Algorithm 1 combines these two conditioning steps, allowing them to be carried out in a single pass. Because DisTaC leverages the already-trained task vectors as the initialization for KD and relies solely on unlabeled data, it incurs minimal computational overhead and imposes only modest practical requirements, yet markedly improves the robustness of existing model merging techniques in challenging scenarios.
Empirically, on eight vision tasks with ViT-B-32/L-14 backbones, DisTaC increased post-merge accuracy by up to 20.8 percentage points and restored the best-performing TSVM merge’s normalized accuracy from 68% to 92% under low-confidence conditions, thereby matching the conventional “ideal” benchmark performance (i.e., merging high-confidence models with uniform task vector norms), all with minimal computational cost. Our contributions are as follows:
@@
  • From our analysis, we present two guidelines for model merging: (i) when the task vector norms differ, it is better to shrink the larger vector rather than stretch the smaller one (Section 6.1); and (ii) when the source models have low confidence, it is more effective to make them overconfident before merging, and then apply a calibration method to the merged model (Section 6.2).
@@ 2 Preliminaries
Let $f(\cdot\,;\bm{\theta}):\mathcal{X}\to\mathbb{R}^{C}$ be a neural network for a $C$-class classification task, parameterized by a vector $\bm{\theta}\in\mathbb{R}^{d}$. The network maps an input vector $\bm{x}\in\mathcal{X}\subseteq\mathbb{R}^{D}$ to a $C$-dimensional logit vector. We target a multi-task scenario comprising $T$ supervised tasks. Let $\bm{\theta}_{\text{pre}}\in\mathbb{R}^{d}$ be the parameters of an open-source pretrained backbone. For each task $t\in\{1,\dots,T\}$, we obtain a model that has already been fine-tuned on the corresponding labeled dataset $\mathcal{D}_{t}=\bigl\{(\bm{x}_{t,i},\bm{y}_{t,i})\bigr\}_{i=1}^{n_{t}},$ yielding task-specific weights $\bm{\theta}_{t}\in\mathbb{R}^{d}$. Each label $\bm{y}_{t,i}\in\{0,1\}^{C}$ is a one-hot vector indicating the ground-truth class.
@@ 2.1 Model Merging for Multi-Task Learning
Recent model merging techniques operate on the task vectors (ilharco2023editingmodelstaskarithmetic) $\bm{\tau}_{t}\;:=\;\bm{\theta}_{t}-\bm{\theta}_{\text{pre}}$ and obtain a single multi–task model by linearly combining them:
$\bm{\theta}_{\text{mtl}}\;=\;\bm{\theta}_{\text{pre}}\;+\;\sum_{t=1}^{T}\bm{P}_{t}\,\bm{\tau}_{t},$ (1)
where each $\bm{P}_{t}\in\mathbb{R}^{d\times d}$ is a method-specific matrix that mitigates inter-task interference.
In the following, we explain the $\bm{P}_{t}$ used in each merging method.
Uniform averaging: $\bm{P}_{t}=\tfrac{1}{T}\,\bm{I}_{d}$.
Task arithmetic (ilharco2023editingmodelstaskarithmetic): $\bm{P}_{t}=\lambda_{t}\,\bm{I}_{d},\quad\text{where}\ \lambda_{t}\in\mathbb{R}$.
Ties-Merging (yadav2023tiesmerging): $\bm{P}_{t}=\lambda_{t}\,\bm{m}_{\text{Ties},t}\,\bm{I}_{d},\quad\text{where}\ \lambda_{t}\in\mathbb{R},\ \bm{m}_{\text{Ties},t}\in\{0,1\}^{d}$. $\bm{m}_{\text{Ties},t}$ is determined by the norm of each weight parameter to mitigate inter-task conflicts.
Consensus Merging (wang2024localizing): $\bm{P}_{t}=\lambda_{t}\,\bm{m}_{\text{Cons},t}\,\bm{I}_{d},\quad\text{where}\ \lambda_{t}\in\mathbb{R},\ \bm{m}_{\text{Cons},t}\in\{0,1\}^{d}$. The framework is the same as Ties-Merging, but the binary mask $\bm{m}_{\text{Cons},t}$ is determined in the following steps. First, create the TALL mask $\bm{m}_{\text{TALL},t}$, which is a binary mask of weights where each element is set to 1 if the norm of $\bm{\tau}_{t}$ is larger than the weighted distance between $\bm{\tau}_{t}$ and $\sum_{t=1}^{T}\bm{\tau}_{t}$. Then, create $\bm{m}_{\text{Cons},t}$, where each element is set to 1 if the corresponding element of $\bm{m}_{\text{TALL},t}$ is 1 in at least $k$ tasks, reflecting agreement among the source models regarding the importance.
TSVM (gargiulo2025task) cannot be expressed within the framework of Eq. (1). Instead, it suppresses task interference by whitening the matrices $\mathbf{U}_{t}$ and $\mathbf{V}_{t}$ obtained from the singular value decomposition of the task vectors $\bm{\tau}_{t}=\mathbf{U}_{t}\mathbf{\Sigma}_{t}\mathbf{V}_{t}^{\top}$.
@@ 2.2 Knowledge Distillation
Knowledge distillation (KD) is a model compression and transfer paradigm in which a compact student network is trained to replicate the behavior of a larger, well-performing teacher network (hinton2015distilling). By minimizing a joint loss that combines ground-truth supervision with a soft-target signal derived from the teacher’s output distribution, the student acquires the teacher’s dark knowledge, namely, fine-grained inter-class relationships encoded in the soft logits, while retaining a substantially smaller parameter footprint. Formally, for a given input $\bm{x}$, let $\bm{z}_{\text{tcr}}:=f(\bm{x}\,;\bm{\theta}_{\text{tcr}})\in\mathbb{R}^{C}$ and $\bm{z}_{\text{stu}}:=f(\bm{x}\,;\bm{\theta}_{\text{stu}})\in\mathbb{R}^{C}$ be the output logits from the teacher and student models, parameterized by $\bm{\theta}_{\text{tcr}}\in\mathbb{R}^{d}$ and $\bm{\theta}_{\text{stu}}\in\mathbb{R}^{d}$, respectively. The KD objective then augments the conventional cross-entropy loss $\mathcal{L}_{\text{CE}}$ with a softened Kullback-Leibler (KL) divergence term:
$\mathcal{L}_{\text{KD}}=(1-\zeta)\,\mathcal{L}_{\text{CE}}\bigl(\bm{z}_{\text{stu}},\,\bm{y}\bigr)\;+\;\zeta\ T_{\text{tcr}}T_{\text{stu}}\,\mathrm{KL}\Bigl(\sigma(\bm{z}_{\text{tcr}}/T_{\text{tcr}})\,\bigl\|\,\sigma(\bm{z}_{\text{stu}}/T_{\text{stu}})\Bigr),$ (2)
@@ 3.1 Task Vector Norm Disparity
To quantify this effect, we fine-tuned CLIP models with Vision Transformers (ViTs) backbones, specifically ViT-B-32, on eight vision tasks as introduced in Section 5.1 with two learning rates, $10^{-5}$ (gray) and $10^{-4}$ (blue), and plotted the resulting task vector norms in Figure 1(b). Across all tasks, we observe a $5$-$7\times$ gap between the two settings. Crucially, the difference is not confined to any particular layer: parameter scales diverge consistently throughout the network, as demonstrated in Section E.1.
Figure 1(a) reports the corresponding merge performance. The gray bars denote the baseline where all eight tasks are fine-tuned with $10^{-5}$, while the blue bars show the average over eight experiments in each of which one task is replaced with a higher learning rate of $10^{-4}$ and the other seven remain unchanged. We measure performance using normalized accuracy. Injecting a single high-norm task vector degrades every merging method, with losses of up to $14\%$. These results confirm that norm discrepancies pose a fundamental obstacle to robust task vector merging.
The detrimental effect of norm disparity on model merging can be explained with a straightforward theoretical analysis formalized as Proposition 1.
Proposition 1.
Let $\bm{\tau}_{1},\bm{\tau}_{2}\in\mathbb{R}^{d}$ with $\|\bm{\tau}_{2}\|>0$, and define $\delta\coloneqq\|\bm{\tau}_{1}\|/\|\bm{\tau}_{2}\|$. Assume $\bm{\tau}_{1}\!\perp\!\bm{\tau}_{2}$. For $\bm{\tau}_{\mathrm{merge}}=\bm{\tau}_{1}+\bm{\tau}_{2}$,
$\cos(\bm{\tau}_{\mathrm{merge}},\bm{\tau}_{2})=\frac{1}{\sqrt{1+\delta^{2}}}\geq 1-\tfrac{1}{2}\delta^{2},\qquad\cos(\bm{\tau}_{\mathrm{merge}},\bm{\tau}_{1})=\frac{\delta}{\sqrt{1+\delta^{2}}}\leq\delta.$
Hence, when $\delta\ll 1$, the merge is nearly perfectly aligned with $\bm{\tau}_{2}$ while its alignment with $\bm{\tau}_{1}$ is at most $O(\delta)$.
Empirically, task vectors are observed to be approximately orthogonal (ilharco2023editingmodelstaskarithmetic); assuming orthogonality, we obtain Proposition 1. The proof is given in Appendix B. This result shows that the merged solution almost entirely inherits the directional characteristics of the high-norm task, while the contribution of the low-norm task vanishes up to $O(\delta)$. Under the Neural Tangent Kernel approximation (jacot2018neural) $\Delta f(x)=f(x;\bm{\theta}_{0}+\bm{\tau}_{\text{merge}})-f(x;\bm{\theta}_{0})\approx\bm{\tau}_{\text{merge}}^{\top}\nabla_{\bm{\theta}}f(x;\bm{\theta}_{0})$, the functional shift from pre-trained model is determined exclusively by the task vector’s direction. Thus, the geometric dominance of the high-norm vector implies that the merged model functionally mimics the high-norm task while failing to preserve the low-norm task’s knowledge, leading to a severe performance drop. Consequently, such norm disparity can cause a severe drop in performance on the low-norm task and thereby degrade the overall effectiveness of the merged model.